# Collaborative Rational Speech Act: Pragmatic Reasoning for Multi-Turn Dialog

### v1

@@ Abstract
As AI systems take on collaborative roles, they must reason about shared goals and beliefs—not just generate fluent language. The Rational Speech Act (RSA) framework offers a principled approach to pragmatic reasoning, but existing extensions face challenges in scaling to multi-turn, collaborative scenarios. In this paper, we introduce Collaborative Rational Speech Act (CRSA), an information-theoretic (IT) extension of RSA that models multi-turn dialog by optimizing a gain function adapted from rate-distortion theory. This gain is an extension of the gain model that is maximized in the original RSA model but takes into account the scenario in which both agents in a conversation have private information and produce utterances conditioned on the dialog. We demonstrate the effectiveness of CRSA on referential games and template-based doctor–patient dialogs in the medical domain. Empirical results show that CRSA yields more consistent, interpretable, and collaborative behavior than existing baselines—paving the way for more pragmatic and socially aware language agents.
@@
Lautaro Estienne1,2, 4, 6 Gabriel Ben Zenou1,4  Nona Naderi2, 4, 6 Jackie Cheung3, 7Pablo Piantanida1,3, 4, 5, 6 1International Laboratory on Learning Systems (ILLS) 2Laboratoire Interdisciplinaire des Sciences du Numérique (LISN) 3Mila - Quebec AI Institute  4CNRS  5CentraleSupélec  6Université Paris-Saclay 7McGill University, Canada CIFAR AI Chair lautaro.estienne@universite-paris-saclay.fr
@@ 1 Introduction
Modeling conversations is central to the development of grounded and useful agentic AI systems, which are increasingly characterized by collaborative interactions between humans and machines. Several applications benefit from dialog systems capable of natural and pragmatic interactions with users. For instance, in the medical domain, conversational agents could support diagnostic interviews (Tu et al., 2025) or serve as tools for physician training in controlled environments (Karunanayake, 2025). In enterprise settings, dialog agents could autonomously handle routine tasks—such as scheduling, data entry, or report generation—freeing human effort for higher-level decision-making (Tupe and Thube, 2025; Satav, 2025). In education, they offer the potential to personalize content delivery, adapting to learners’ styles and paces (Nabhani et al., 2025; Vorobyeva et al., 2025). While such applications are still emerging, a key enabler is the development of models that can manage collaborative, goal-oriented interactions in a robust and interpretable manner.
To succeed in real-world settings, dialog generative-based models must do more than generate fluent language—they must track shared tasks to communicate meaningfully and contextually (Lin et al., 2024). For example, a physician in a diagnostic exchange refines hypotheses as the conversation evolves, requiring interpretable and scalable frameworks for reliable interaction.
Yet, many existing models prioritize task-specific response generation (He et al., 2017; Jiang et al., 2019; Meta Fundamental AI Research Diplomacy Team et al.(2022)Meta Fundamental AI Research Diplomacy Team (FAIR), Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, Jacob, Komeili, Konath, Kwon, Lerer, Lewis, Miller, Mitts, Renduchintala, Roller, Rowe, Shi, Spisak, Wei, Wu, Zhang, and Zijlstra, FAIR), or optimize for superficial conversation properties using narrowly defined objectives (Khani et al., 2018; Dafoe et al., 2020; Lin et al., 2024; Jeon et al., 2020). While these methods often yield strong performance, they typically lack principled foundations and depend on ad hoc design choices.
@@
Contributions
@@ 2 Related work
The Rational Speech Act (RSA) framework (Frank and Goodman, 2012) serves as a model for pragmatic communication designed to emulate human behavior in linguistic tasks Degen et al. (2020); Bergen et al. (2016); Herbstritt and Franke (2019). This framework is both conceptually intuitive and computationally versatile, making it readily adaptable for integration with neural language models to tackle more intricate challenges, including machine translation (Cohn-Gordon and Goodman, 2019), image captioning (Cohn-Gordon et al., 2018), controllable text generation (Shen et al., 2019; Wang and Demberg, 2024). Extensions to the original RSA framework have been proposed to accommodate more complex scenarios. For instance, adaptations have addressed cases where agents lack shared vocabularies (Bergen et al., 2016) or where common ground evolves dynamically during interaction (Degen et al., 2015). A comprehensive overview of RSA’s development and its numerous variants is provided by Degen (2023).
@@
Multiple works frame a collaborative or task-oriented dialog as a Partially Observable Markov Decision Process (POMDP) (Williams and Young, 2007), which provide a suitable framework to model end-to-end networks on specific tasks (Wen et al., 2017; Jiang et al., 2019). Reinforcement Learning has been widely used in this context, in order to provide interpretable and trackable training procedures that incorporate the structure of the dialog in their policy training or decoding strategy (Lin et al., 2024; Li et al., 2016; Xu et al., 2025). Related to this, game-theoretic perspective has also been used in dialog modeling (Jeon et al., 2020; Lin et al., 2022). In this context, multiple tasks and datasets have been developed to evaluate dialog modeling (He et al., 2017; Khani et al., 2018; Macherla et al., 2023), usually by assessing the task performance and the similarity with human conversations. The RSA model has also found applications in dialog systems, often complementing neural models to enhance agent self-awareness (Kim et al., 2020) or to improve the interpretation of emotional subtext (Kim et al., 2021).
@@ 3 Review of the RSA Model from the Lens of Information Theory
Figure 1: RSA variants proposed in this work(1(b), 1(c)) compared to the original one (1(a)).
Fig. 1(a) presents a schematic view of the classic RSA model from an information-theoretic perspective. Here, a meaning $m\in\mathcal{M}$ is received by the speaker $S:\mathcal{M}\times\mathcal{U}\rightarrow[0,1]$ who uses it to produce a posterior probability $S(u|m)$ for all possible utterances $u\in\mathcal{U}$. The utterance $u$ is then transmitted to the listener $L:\mathcal{U}\times\mathcal{M}\rightarrow[0,1]$ who produces a posterior $L(m|u)$ for all possible reconstructions $m\in\mathcal{M}$ of the meaning $m$ that the speaker is trying to convey. Additionally, there is a distribution $P:\mathcal{M}\rightarrow[0,1]$ that is known by the two agents and represents the prior of the meanings. Finally, the function $C:\mathcal{U}\rightarrow\mathbb{R}$ assigns a prior cost value to each utterance produced by the speaker.
@@
Zaslavsky et al. (2021) show that this iteration process is equivalent to maximize the next objective:
@@ 4.1 Modeling private meanings (YRSA)
To extend the RSA model to bidirectional dialogue with explicit task modeling, we first distinguish between private meanings and shared task outcomes. In real conversations, each participant holds their own prior knowledge and worldview, which may differ from that of their interlocutor. In our example of a dialogue between a patient and a physician: the patient must describe their symptoms, which are not directly observable by the physician, while the physician brings medical expertise the patient lacks. Both types of knowledge are essential to determine the appropriate diagnosis or treatment plan. Notably, neither the patient’s symptoms nor the physician’s prior knowledge fully align with the shared goal of the conversation, i.e. the identification of a suitable medical outcome.
In this context, we identify the need of representing a private set of meanings $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$ for each agent, which may or may not match. In addition, the result $y$ of the shared task is going to be represented with a separate space $\mathcal{Y}$ that contains all the possible outcomes of it. For simplicity, we will assume that all these are discrete spaces. Fig. 1(b) represents a schematic of this model. We will refer to this extension as the YRSA model.
The YRSA model redefines the notion of prior from the classic RSA framework by conditioning the dialogue on the joint realization of the agents’ private meanings ($m_{A}$, $m_{B}$) and the shared task target $y$, which together define the context in which the interaction unfolds. Importantly, we assume for the development of our model that both the realizations and the joint distribution of these three variables do not change over time during the conversation. This implies that the prior is completely defined by the joint distribution $P:\mathcal{M}_{A}\times\mathcal{M}_{B}\times\mathcal{Y}\rightarrow[0,1]$ given to both agents.
@@ 4.2 The CRSA Model
Effective collaboration requires not only modeling agents’ private meanings and the shared task, but also supporting multi-turn dialogue. In a medical consultation, for instance, the patient shares symptoms and background, while the physician asks questions, proposes diagnoses, and recommends treatments. To capture such interactions, we denote the speaker’s utterance at turn $t$ as $U_{t}$, and the dialogue history up to that point as $W_{t}=(U_{1},\ldots,U_{t-1})$, representing the sequence of prior exchanges.
4.2.1 Modeling multi-turn dialog with explicit history (baseline)
The attempt of previous approaches to incorporate the history of the conversation to the RSA model relies on defining the lexicon (or directly the literal listener/speaker) as a function of each turn (Wang and Demberg, 2024; Kim et al., 2020; Lin et al., 2022). In many cases, this lexicon is given by the output of a neural language model and can be very robust to the evolving dialog. However, that variant of the RSA does not correspond to maximizing the gain of Eq. (1), but a modified version of it in which $U_{t}$ is replaced by $(U_{t},W_{t})$:
@@ 4.2.1 Modeling multi-turn dialog with explicit history (baseline)
The issue with Eq. (3) is that the speaker’s utterance $U_{t}$ at turn $t$ is modeled jointly with the dialogue history $W_{t}$, rather than being explicitly conditioned on it. To express the gain in terms of the conditional entropy of the current utterance alone, we condition it on both the dialogue history $W_{t}$ and the speaker’s intended meaning $M$, rather than on $M$ alone. In Section 4.2.2, we formally introduce the corresponding expressions of the CRSA model, which incorporates this notion of multi-turn conditioned to the past utterances, as well as private meanings and target task.
@@ 4.2.2 Equations of the CRSA model
Figure 1(c) illustrates our extension of the YRSA model to the collaborative setting. As in the original setup, agents alternate roles—one acting as the speaker, the other as the listener—to achieve a shared task. Each agent has access to a private meaning space, $\mathcal{M}_{A}$ or $\mathcal{M}_{B}$, which remains hidden from their counterpart. Then, at turn $t$, the private meanings of the speaker will correspond to the meanings of the agent playing the role of the speaker and vice-versa. We refer as $\mathcal{M}_{S_{t}}$ and $\mathcal{M}_{L_{t}}$ to the private meanings of the speaker and the listener at turn $t$, respectively. Both agents also have access to the conversation history, denoted as $w_{t}=(u_{1},\ldots,u_{t-1})\in\mathcal{W}_{t}\triangleq\mathcal{U}_{1}\times\cdots\times\mathcal{U}_{t-1}$, where each $\mathcal{U}_{i}$ represents the space of possible utterances at turn $i$. The shared objective is to jointly predict a target class $y$ from a finite discrete set $\mathcal{Y}$.
@@
Formally, we define the distribution of each agent at turn $t$. The speaker $S_{t}:\mathcal{M}_{S_{t}}\times\mathcal{U}_{t}\times\mathcal{W}_{t}\rightarrow[0,1]$ produces a posterior $S_{t}(u_{t}|m_{S_{t}},w_{t})$ that depends on its private meaning $m_{S_{t}}$ and the past utterances $w_{t}$. On the other hand, the listener $L_{t}:\mathcal{M}_{L_{t}}\times\mathcal{U}\times\mathcal{W}_{t}\times\mathcal{Y}\rightarrow[0,1]$ is represented by the posterior $L_{t}(y|m_{L_{t}},u_{t},w_{t})$ which is independent of the private meanings of the speaker.
Building on the gain function in Eq. (1), we extend the joint speaker distribution and listener utility to incorporate private meanings and multi-turn dialogue:
@@
where we replace $B^{\prime}_{t}(m_{S},m_{L},y)=$
@@
Regarding algorithmic complexity, we find that at turn $t$, these new set of equations scale as $\mathcal{O}\left(K\cdot|\mathcal{M}_{A}|\cdot|\mathcal{M}_{B}|\cdot|\mathcal{Y}|\cdot|\mathcal{U}_{t}|\right)$ where $K$ is the number of iterations to produce the pragmatic agents. In contrast, the classic RSA equations scale as $\mathcal{O}\left(K\cdot|\mathcal{M}|\cdot|\mathcal{U}|\right)$.
@@ 5.1 Experimental set-up
For this simulation we consider that the set $\mathcal{U}_{t}$ of possible utterances at turn $t$ is always ($\forall t$) the set $\mathcal{U}_{t}=\{1,\ldots,N\}$ representing the messages of the form “A1 card may be at position $n$ with $n\in\mathcal{U}_{t}$. For the set $\mathcal{Y}$ of possible classes, the results can be as well “A1 card may be at position $n$, with the addition that there is also the possibility of “There is no A1 card”. That is, $\mathcal{Y}=\{0,1,\ldots,N$} with 00 representing the mentioned possibility. Regarding the meaning spaces, they correspond to the possible sequences of length $N$ that can be obtained combining without replacement the letters A and B (for agent A) and the numbers 1 and 2 (for agent B). That is, for instance if $N=3$, $\mathcal{M}_{A}=\{\textrm{AAA},\textrm{AAB},\ldots,\textrm{BBB}\}$ and $\mathcal{M}_{B}=\{111,112,\ldots,222\}$. Finally, the prior distribution $P(m_{A},m_{B},y)$ can be defined as follows:
@@
Since this is a reference game, we adopt the listener’s perspective. In all cases, the literal listener is initialized using Eq. (4.2.2), and different model variants are defined based on the update equations and the specification of the lexicon $\mathcal{L}_{u_{t},w_{t}}(m_{S_{t}})$.
@@
    $\mathcal{L}(u_{t},m_{S_{t}})=\begin{cases}1&\text{if $m_{S_{t}}$ contains A/1 at}\\[-5.0pt] &\text{position $n$ and $u_{t}=n$}\\ 1&\text{if there is no A/1 in $m_{S_{t}}$}\\ 0&\mbox{otherwise}\end{cases}$ (9)
@@
    We expect efficient conversational behavior in this game to involve repeating an utterance only to confirm the correct A1 card position. If the correct position is identified, agents should repeat the utterance until the round ends; otherwise, repeating it would be inefficient. The rule in Eq. (10) explicitly encodes this behavior.
  • YRSA: We initialize the listener using the YRSA iterative equations and the lexicon from Eq. (9), effectively applying the RSA iteration in the setting where each agent holds a private meaning—that is, the standard YRSA setup.
  • YRSA-$W_{t}$: The same as the one above but using Eq. (10) as lexicon instead of Eq. (9).
  • Literal: In this case, there is no iteration and we simply use Eq. (4.2.2) to predict the target. We use lexicon of Eq (9).
  • Literal-$W_{t}$: This is the same as above but using Eq. (10) as lexicon.

### v2

@@ Abstract
As AI systems take on collaborative roles, they must reason about shared goals and beliefs—not just generate fluent language. The Rational Speech Act (RSA) framework offers a principled approach to pragmatic reasoning, but existing extensions face challenges in scaling to multi-turn, collaborative scenarios. In this paper, we introduce Collaborative Rational Speech Act (CRSA), an information-theoretic (IT) extension of RSA that models multi-turn dialog by optimizing a gain function adapted from rate-distortion theory. This gain is an extension of the gain model that is maximized in the original RSA model but takes into account the scenario in which both agents in a conversation have private information and produce utterances conditioned on the dialog. We demonstrate the effectiveness of CRSA on referential games and template-based doctor-patient dialogs in the medical domain. Empirical results show that CRSA yields more consistent, interpretable, and collaborative behavior than existing baselines, paving the way for more pragmatically competent language agents.
@@
Lautaro Estienne1,2, 4, 6  Gabriel Ben Zenou1,4   Nona Naderi2, 4, 6 Jackie Chi Kit Cheung3, 7Pablo Piantanida1,3, 4, 5, 6 1International Laboratory on Learning Systems (ILLS) 2Laboratoire Interdisciplinaire des Sciences du Numérique (LISN) 3Mila - Quebec AI Institute  4CNRS  5CentraleSupélec  6Université Paris-Saclay 7McGill University, Canada CIFAR AI Chair lautaro.estienne@universite-paris-saclay.fr
@@ 1 Introduction
Modeling conversations is central to the development of grounded and useful agentic AI systems, which are increasingly characterized by collaborative interactions between humans and machines. Several applications benefit from dialog systems capable of natural interactions with users. For instance, in the medical domain, conversational agents could support diagnostic interviews (Tu et al., 2025) or serve as tools for physician training in controlled environments (Karunanayake, 2025). In enterprise settings, dialog agents could autonomously handle routine tasks—such as scheduling, data entry, or report generation—freeing human effort for higher-level decision-making (Tupe and Thube, 2025; Satav, 2025). In education, they offer the potential to personalize content delivery, adapting to learners’ styles and paces (Nabhani et al., 2025; Vorobyeva et al., 2025). While such applications are still emerging, a key enabler is the development of models that can manage collaborative, goal-oriented interactions in a robust and interpretable manner.
To succeed in real-world settings, dialog generative-based models must do more than generate fluent language—they must track shared tasks to communicate meaningfully in context (Lin et al., 2024). For example, a physician in a diagnostic exchange refines hypotheses as the conversation evolves, requiring interpretable and scalable frameworks for reliable interaction.
Yet, many existing models prioritize task-specific response generation (He et al., 2017; Jiang et al., 2019; Meta Fundamental AI Research Diplomacy Team et al.(2022)Meta Fundamental AI Research Diplomacy Team (FAIR), Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, Jacob, Komeili, Konath, Kwon, Lerer, Lewis, Miller, Mitts, Renduchintala, Roller, Rowe, Shi, Spisak, Wei, Wu, Zhang, and Zijlstra, FAIR), or optimize for superficial conversation properties using narrowly defined objectives (Khani et al., 2018; Dafoe et al., 2020; Lin et al., 2024; Jeon et al., 2020). While these methods often yield strong performance, they typically lack principled foundations, leading to task-specific solutions that struggle to generalize or remain robust under shifting conditions.
@@
In this paper, we introduce the Collaborative Rational Speech Act (CRSA), an IT grounded extension of the RSA framework designed to model multi-turn, collaborative dialogs. CRSA optimizes a gain function that is a generalization of the one proposed by Zaslavsky et al. (2021) for the multiple-turns scenario. The resulting model provides a tool to model an estimation of the target of the joint task and the belief that each agent has on their in interlocutor’s private information, and it can be used with large language models (LLMs). We evaluate CRSA in referential game settings and semi-automatic generated conversations between doctor and patients for extracting a medical diagnosis. [^1] Code is available at https://github.com/LautaroEst/crsa
@@ 2 Related work
The Rational Speech Act (RSA) framework (Frank and Goodman, 2012) serves as a model for pragmatic communication designed to emulate human behavior in linguistic tasks Degen et al. (2020); Bergen et al. (2016); Herbstritt and Franke (2019); Spinoso-Di Piano et al. (2025). This framework is both conceptually intuitive and computationally versatile, making it readily adaptable for integration with neural language models to tackle more intricate challenges, including machine translation (Cohn-Gordon and Goodman, 2019), image captioning (Cohn-Gordon et al., 2018), controllable text generation (Shen et al., 2019; Wang and Demberg, 2024; Darrin et al., 2024). Extensions to the original RSA framework have been proposed to accommodate more complex scenarios. For instance, adaptations have addressed cases where agents lack shared vocabularies (Bergen et al., 2016) or where common ground evolves dynamically during interaction (Degen et al., 2015). A comprehensive overview of RSA’s development and its numerous variants is provided by Degen (2023).
@@
Multiple works frame a collaborative or task-oriented dialog as a Partially Observable Markov Decision Process (POMDP) (Williams and Young, 2007), which provide a suitable framework to model end-to-end networks on specific tasks (Wen et al., 2017; Jiang et al., 2019). Reinforcement learning has been widely used in this context in order to provide interpretable and trackable training procedures that incorporate the structure of the dialog in their policy training or decoding strategy (Lin et al., 2024; Li et al., 2016; Xu et al., 2025). Related to this, game-theoretic perspective has also been used in dialog modeling (Jeon et al., 2020; Lin et al., 2022). In this context, multiple tasks and datasets have been developed to evaluate dialog modeling (He et al., 2017; Khani et al., 2018; Macherla et al., 2023), usually by assessing the task performance and the similarity with human conversations. The RSA model has also found applications in dialog systems, often complementing neural models to enhance agent consistency given persona (Kim et al., 2020) or to improve the interpretation of emotional subtext (Kim et al., 2021).
@@ 3 Review of the RSA Model from the Lens of Information Theory
Figure 1: RSA variants proposed in this work (YRSA 1(b), CRSA 1(c)) compared to the original one (RSA 1(a)).
Figure 1(a) presents a schematic view of the classic RSA model from an information-theoretic perspective. Here, a meaning $m\in\mathcal{M}$ is received by the speaker $S:\mathcal{M}\times\mathcal{U}\rightarrow[0,1]$ who uses it to produce a posterior probability $S(u|m)$ for all possible utterances $u\in\mathcal{U}$. The utterance $u$ is then transmitted to the listener $L:\mathcal{U}\times\mathcal{M}\rightarrow[0,1]$ who produces a posterior $L(m|u)$ for all possible reconstructions $m\in\mathcal{M}$ of the meaning $m$ that the speaker is trying to convey. Additionally, there is a distribution $P:\mathcal{M}\rightarrow[0,1]$ that is known by the two agents and represents the prior of the meanings. Finally, the function $C:\mathcal{U}\rightarrow\mathbb{R}$ assigns a prior cost value to each utterance produced by the speaker. [^2] Note that the prior and cost functions are not shown in Figure 1 for clarity.
@@
Zaslavsky et al. (2021) show that this iteration process is equivalent to maximizing the following objective:
@@ 4.1 Modeling private meanings (YRSA)
To extend the RSA model to bidirectional dialog with explicit task modeling, we first distinguish between private meanings and shared task outcomes. In real conversations, each participant holds their own prior knowledge and worldview, which may differ from that of their interlocutor. In our example of a dialog between a patient and a physician: the patient must describe their symptoms, which are not directly observable by the physician, while the physician brings medical expertise the patient lacks. Both types of knowledge are essential for determining the appropriate diagnosis or treatment plan. Notably, neither the patient’s symptoms nor the physician’s prior knowledge fully capture the shared goal of the conversation, i.e. the identification of a suitable medical outcome.
In this context, we identify the need of representing a private set of meanings $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$ for each agent, which may or may not match. In addition, the result $y$ of the shared task is going to be represented with a separate space $\mathcal{Y}$ that contains all the possible outcomes of it. For simplicity, we will assume that all these are discrete spaces. Figure 1(b) represents a schematic of this model. We will refer to this extension as the YRSA model.
The YRSA model redefines the notion of prior from the classic RSA framework by conditioning the dialog on the joint realization of the agents’ private meanings ($m_{A}$, $m_{B}$) and the shared task target $y$, which together define the context in which the interaction unfolds. Importantly, we assume for the development of our model that both the realizations and the joint distribution of these three variables do not change over time during the conversation. This implies that the prior is completely defined by the joint distribution $P:\mathcal{M}_{A}\times\mathcal{M}_{B}\times\mathcal{Y}\rightarrow[0,1]$ given to both agents.
@@ 4.2 The CRSA Model
Effective collaboration requires not only modeling agents’ private meanings and the shared task, but also supporting multi-turn dialog. In a medical consultation, for instance, the patient shares symptoms and background, while the physician asks questions, proposes diagnoses, and recommends treatments. To capture such interactions, we denote the speaker’s utterance at turn $t$ as $U_{t}$, and the dialog history up to that point as $W_{t}=(U_{1},\ldots,U_{t-1})$, representing the sequence of prior exchanges.
The attempt of previous approaches to incorporate the history of the conversation to the RSA model relies on defining the lexicon (or directly the literal listener/speaker) as a function of each turn (Wang and Demberg, 2024; Kim et al., 2020; Lin et al., 2022). In many cases, this lexicon is given by the output of a neural language model and can be very robust to the evolving dialog. However, that variant of the RSA does not correspond to maximizing the gain of Equation (1), but a modified version of it in which $U_{t}$ is replaced by $(U_{t},W_{t})$:
@@
The issue with Equation (3) is that the speaker’s utterance $U_{t}$ at turn $t$ is modeled jointly with the dialog history $W_{t}$, rather than being explicitly conditioned on it. To express the gain in terms of the conditional entropy of the current utterance alone, we condition it on both the dialog history $W_{t}$ and the speaker’s intended meaning $M$, rather than on $M$ alone. In Section 4.2.1, we formally introduce the corresponding expressions of the CRSA model, which incorporates this notion of multi-turn conditioned to the past utterances, as well as private meanings and target task.
@@ 4.2.1 Equations of the CRSA model
Figure 1(c) illustrates our extension of the YRSA model to the collaborative setting. As in the original setup, agents alternate roles—one acting as the speaker, the other as the listener—to achieve a shared task. Each agent has access to a private meaning space, $\mathcal{M}_{A}$ or $\mathcal{M}_{B}$, which remains hidden from their counterpart. Then, at turn $t$, the private meanings of the speaker will correspond to the meanings of the agent playing the role of the speaker and vice-versa. We refer as $\mathcal{M}_{S_{t}}$ and $\mathcal{M}_{L_{t}}$ to the private meanings of the speaker and the listener at turn $t$, respectively. For instance, if speaker A starts the conversation, $\mathcal{M}_{S_{1}}=\mathcal{M}_{A}$ and $\mathcal{M}_{L_{1}}=\mathcal{M}_{B}$. Both agents also have access to the conversation history, denoted as $w_{t}=(u_{1},\ldots,u_{t-1})\in\mathcal{W}_{t}\triangleq\mathcal{U}_{1}\times\cdots\times\mathcal{U}_{t-1}$, where each $\mathcal{U}_{i}$ represents the space of possible utterances at turn $i$. The shared objective is to jointly predict a target class $y$ from a finite discrete set $\mathcal{Y}$.
@@
Formally, we define the distribution of each agent at turn $t$. The speaker $S_{t}:\mathcal{M}_{S_{t}}\times\mathcal{U}_{t}\times\mathcal{W}_{t}\rightarrow[0,1]$ produces a posterior $S_{t}(u_{t}|m_{S_{t}},w_{t})$ that depends on its private meaning $m_{S_{t}}$ and the past utterances $w_{t}$. On the other hand, the listener $L_{t}:\mathcal{M}_{L_{t}}\times\mathcal{U}_{t}\times\mathcal{W}_{t}\times\mathcal{Y}\rightarrow[0,1]$ is represented by the posterior $L_{t}(y|m_{L_{t}},u_{t},w_{t})$ which is independent of the private meanings of the speaker.
Building on the gain function in Equation (1), we extend the joint speaker distribution and listener utility to incorporate private meanings and multi-turn dialog:
@@
where we replace [^4] For simplify notation, we removed the $t$ subindex. $B^{\prime}_{t}(m_{S},m_{L},y)=$
@@
Algorithmic complexity.
We find that at turn $t$, these new set of equations scale as $\mathcal{O}\left(K\cdot|\mathcal{M}_{A}|\cdot|\mathcal{M}_{B}|\cdot|\mathcal{Y}|\cdot|\mathcal{U}_{t}|\right)$ where $K$ is the number of iterations to produce the pragmatic agents. In contrast, the classic RSA equations scale as $\mathcal{O}\left(K\cdot|\mathcal{M}|\cdot|\mathcal{U}|\right)$.
@@ 5.1 Experimental set-up
For this simulation, we consider that the set $\mathcal{U}_{t}$ of possible utterances at turn $t$ is always ($\forall t$) the set $\mathcal{U}_{t}=\{1,\ldots,N\}$ representing the messages of the form “A1 card may be at position $n$ with $n\in\mathcal{U}_{t}$. For the set $\mathcal{Y}$ of possible classes, the results can be as well “A1 card may be at position $n$, with the addition that there is also the possibility of “There is no A1 card”. That is, $\mathcal{Y}=\{0,1,\ldots,N$} with 0 representing the mentioned possibility. Regarding the meaning spaces, they correspond to the possible sequences of length $N$ that can be obtained combining without replacement the letters A and B (for agent A) and the numbers 1 and 2 (for agent B). That is, for instance if $N=3$, $\mathcal{M}_{A}=\{\textrm{AAA},\textrm{AAB},\ldots,\textrm{BBB}\}$ and $\mathcal{M}_{B}=\{111,112,\ldots,222\}$. Finally, the prior distribution $P(m_{A},m_{B},y)$ can be defined as follows:
@@
Since this is a reference game, we adopt the listener’s perspective. In all cases, the literal listener is initialized using Equation (4.2.1), and different model variants are defined based on the update equations and the specification of the lexicon $\mathcal{L}_{u_{t},w_{t}}(m_{S_{t}})$.
@@
[…]
@@
    We expect efficient conversational behavior in this game to involve repeating an utterance only to confirm the correct A1 card position. If the correct position is identified, agents should repeat the utterance until the round ends; otherwise, repeating it would be inefficient. The rule in Equation (10) explicitly encodes this behavior.
  • YRSA: We initialize the listener using the YRSA iterative equations and the lexicon from Equation (9), effectively applying the RSA iteration in the setting where each agent holds a private meaning—that is, the standard YRSA setup.
  • YRSA-$W_{t}$: The same as the one above but using Equation (10) as lexicon instead of Equation (9).
  • Literal: In this case, there is no iteration and we simply use Equation (4.2.1) to predict the target. We use lexicon of Equation (9).
  • Literal-$W_{t}$: This is the same as above but using Equation (10) as lexicon.